A Bibliography of the Proceedings Volumes of the IEEE Symposia on the Foundations of Computer Science

(1975–date)

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

14 October 2017 Version 2.06

Title word cross-reference

13th [13]. 14th [14]. 15th [15]. 16th [16]. 17th [17]. 18th [18]. 19th [19]. 1st [1].

2002 [43]. 20th [20]. 21st [21]. 22nd [22]. 23rd [23]. 24rd [24]. 25th [25]. 26-28 [7]. 26th [26]. 27th [27]. 28th [28]. 29th [29].

30th [30]. 31st [31]. 32nd [32]. 33rd [33]. 34th [34]. 35th [35]. 36th [36]. 37th [37]. 38th [38]. 39th [39].

40th [40]. 41st [41]. 42nd [42]. 43rd [43]. 44th [44]. 45th [45]. 46th [46]. 47th [47].

48th [48]. 49th [49].

50th [50]. 51st [51]. 52nd [52]. 53rd [53]. 54th [54].

Alto [34, 39]. Angeles [28]. Ann [6, 19]. Annual [35, 1, 3, 8, 11, 13, 14, 15, 16, 20, 22, 24, 31, 33, 34, 36, 37, 38, 39, 40, 41, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 4, 5, 6, 7, 9, 10, 12, 17, 18, 19, 21, 23, 25, 26, 27, 28, 29, 30, 32]. Arbor [6, 19]. Arizona [24]. Atlanta [50]. Automata [8, 9, 10, 11, 12, 13, 14, 15, 19, 7].

BC [43]. Beach [38, 41]. Berkeley [7, 16, 47, 54]. Brunswick [53]. Burlington [37].

CA [54]. Calif [7, 11]. California [7, 16, 28, 34, 39, 41, 52, 47]. called [19]. Cambridge [44]. Canada [10, 27, 43]. Carolina [30]. Chicago [3, 4, 23]. Circuit [1, 2, 3, 4, 5, 6]. City [40]. Computer [35, 14, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54]. Conference [12, 6, 7, 8, 9, 10, 11].

Department [14]. Design [1, 2, 3, 4, 5, 6]. Detroit [2].

East [12]. Eighth [8]. Eleventh [11].

Fe [35]. fifth [5]. Florida [25, 38]. FOCS [43, 44, 45, 46, 47, 48]. formerly [19]. Foundations [35, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54]. fourth [4].

Georgia [50].

Houston [17]. Hyatt [53].

IEEE [43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54]. Ill [4]. Illinois [23]. Iowa [14]. Island [18, 25, 48]. Italy [45].

Jersey [53]. Juan [20, 32].

Lansing [12]. Logical [1, 2, 3, 4, 5, 6]. Louis [31].

Massachusetts [44]. Mexico [35]. MI [2]. Miami [38]. Mich [6, 12]. Michigan [6, 19]. Milwaukee [36]. Missouri [31]. Monica N.J. [5]. Nashville [22]. Nevada [42, 51]. ninth [9]. North [30]. November [35, 5, 18, 23, 24, 30, 34, 39, 41, 43].

Oct [19, 20]. October [1, 2, 3, 4, 6, 7, 8, 9, 10, 11, 12, 13, 14, 16, 17, 18, 21, 22, 25, 26, 27, 28, 29, 30, 31, 32, 33, 36, 37, 38, 40, 42, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54]. Ontario [10]. Oregon [26]. Orleans [15].

Palm [52]. Palo [34, 39]. papers [1, 6, 7, 8, 9, 10, 11, 12, 22, 33, 34]. Park [30]. Pennsylvania [33, 46, 49]. Philadelphia [49]. Pittsburgh [33, 46]. Plains [29]. Portland [26]. presented [6, 7, 8, 9, 10, 11, 12]. Princeton [5]. Proceedings [35, 3, 31, 32, 4, 5, 33, 34, 39, 41, 42, 43, 44, 45, 48, 49, 50, 51, 52, 53, 54]. program [1]. Providence [18, 48]. Puerto [20, 32].

record [6, 7, 8, 9, 10, 11, 12]. Redondo [41]. Regency [53]. Research [30]. Rhode [18, 48]. Rico [20, 32]. Rome [45].

San [20, 32]. Santa [35, 11]. Schenectady [9]. Science [35, 14, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54]. Selected [1]. seventh [7]. Singer [25]. sixth [6]. Springs [52]. St [31]. Switching [1, 2, 3, 4, 5, 8, 9, 10, 11, 12, 13, 14, 15, 19, 6, 7]. Symposium [35, 1, 3, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 4, 5, 6, 7]. Syracuse [21].

Tennessee [22]. tenth [10]. Texas [8, 17]. Theory [1, 2, 3, 8, 9, 10, 11, 12, 13, 14, 15, 19, 4, 5, 6, 7]. Third [3]. Toronto [27]. Triangle [30].

Tucson [24]. twelfth [12].

University [6, 7, 8, 14, 15]. USA [42, 46, 49, 50, 51, 52, 53, 54].

Vancouver [43]. Vegas [42, 51]. Vermont [37].

Waterloo [10]. White [29]. Wisconsin [36].

York [9, 21, 29, 40].

References

IEEE:1960:SCT

[1] IEEE, editor. Switching Circuit Theory and Logical Design. 1st Annual Symposium. Selected papers and program. October 9–14, 1960. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1960. CO-DEN ASFPDV. ISSN 0272-5428. LCCN ????

IEEE:1961:SCT

[2] IEEE, editor. Switching Circuit Theory and Logical Design. October 17–20, 1961, Detroit, MI. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1961. CO-DEN ASFPDV. ISSN 0272-5428. LCCN ????

IEEE:1962:PTA

[3] IEEE, editor. Proceedings of the Third Annual Symposium on Switching Circuit Theory and Logical Design: Chicago, October 7–12, 1962. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1962. CO-DEN ASFPDV. ISSN 0272-5428. LCCN ????

IEEE:1963:SCT

[4] IEEE, editor. Switching circuit theory and logical design: proceedings of the fourth annual symposium: Chicago, Ill., October 28–30, 1963. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1963. CO-DEN ASFPDV. ISSN 0272-5428. LCCN TK7885.A1 S92 1963-64.

IEEE:1964:SCT

[5] IEEE, editor. Switching circuit theory and logical design: proceedings of the fifth annual symposium: Princeton, N.J., November 11–13, 1964. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1964. CODEN ASFPDV. ISSN 0272-5428. LCCN TK7885.A1 S92 1963-64.

IEEE:1965:ICR

[6] IEEE, editor. 1965 IEEE conference record on switching circuit theory and logical design; papers presented at the sixth annual symposium, University of Michigan, Ann Arbor, Mich., October 6– 8, 1965. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1965. CO-DEN ASFPDV. ISSN 0272-5428. LCCN QA267 .S95 1965.

IEEE:1966:ICR

[7] IEEE, editor. IEEE conference record of 1966 seventh annual symposium on switching and automata theory : papers presented at the seventh annual symposium : University of California, Berkeley, Calif., October 26-28, 1966. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1966. CODEN ASFPDV. ISSN 0272-5428. LCCN TK7885.A1 S92 1966.

IEEE:1967:ICR

[8] IEEE, editor. IEEE conference record of 1967 Eighth Annual Symposium on Switching and Automata Theory : papers presented at the Eighth Annual Symposium, University of Texas, October 18– 20, 1967. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1967. CODEN ASFPDV. ISSN 0272-5428. LCCN ????

IEEE:1968:ICR

[9] IEEE, editor. IEEE conference record of 1968 ninth annual Symposium on Switching and Automata Theory : papers presented at the ninth annual Symposium, Schenectady, New York, October 15–18, 1968. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1968. CO-DEN ASFPDV. ISSN 0272-5428. LCCN TK7885.A1 S92 1968.

IEEE:1969:ICR

[10] IEEE, editor. IEEE conference record of 1969 tenth annual Symposium on Switching and Automata Theory : papers presented at the tenth annual Symposium, Waterloo, Ontario, Canada, October 15– 17, 1969. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1969. CO-DEN ASFPDV. ISSN 0272-5428. LCCN TK7885.A1 S92 1969.

IEEE:1970:ICR

[11] IEEE, editor. IEEE conference record of 1970 Eleventh Annual Symposium on Switching and Automata Theory : papers presented at the Eleventh Annual Symposium, Santa Monica, Calif., October 28–30, 1970. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1970. CO-DEN ASFPDV. ISSN 0272-5428. LCCN TK7885.A1 S92 1970.

IEEE:1971:CRT

[12] IEEE, editor. Conference record, 1971 twelfth annual Symposium on Switching and Automata Theory : papers presented at the twelfth annual symposium, East Lansing, Mich., October 13–15, 1971. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1971. CO-DEN ASFPDV. ISSN 0272-5428. LCCN TK7885.A1 S92 1971.

IEEE:1972:ASS

[13] IEEE, editor. 13th Annual Symposium on Switching and Automata Theory, October 25–27, 1972. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1972. CO-DEN ASFPDV. ISSN 0272-5428. LCCN TK7885.A1 S92 1972.

IEEE:1973:ASS

[14] IEEE, editor. 14th Annual Symposium on Switching and Automata Theory: Computer Science Department of the University of Iowa, October 15–17, 1973. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1973. CODEN ASFPDV. ISSN 0272-5428. LCCN QA268.5 .S9 1973.

IEEE:1974:ASS

[15] IEEE, editor. 15th Annual Symposium on Switching and Automata Theory: University of New Orleans. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1974. CO-DEN ASFPDV. ISSN 0272-5428. LCCN TK7885.A1 S92 1974.

IEEE:1975:ASF

[16] IEEE, editor. 16th Annual Symposium on Foundations of Computer Science, October 13–15, 1975, Berkeley, California. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1975. CODEN ASF-PDV. ISBN ???? ISSN 0272-5428. LCCN QA 267 S95 1975.

IEEE:1976:ASF

[17] IEEE, editor. 17th annual Symposium on Foundations of Computer Science, October 25–27, 1976, Houston, Texas. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1976. CODEN ASFPDV. ISBN ???? ISSN 0272-5428. LCCN QA75.5 .S97 1976.

IEEE:1977:ASF

[18] IEEE, editor. 18th annual Symposium on Foundations of Computer Science, October 31–November 2, 1977, Providence, Rhode Island. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1977. CO-DEN ASFPDV. ISBN ???? ISSN 0272-5428. LCCN TK7885.A1 S92 1977.

IEEE:1978:ASF

[19] IEEE, editor. 19th annual Symposium on Foundations of Computer Science (formerly called the Annual Symposium on Switching and Automata Theory) Oct. 16–18, 1978, Ann Arbor, Michigan. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1978. CODEN ASFPDV. ISBN ???? ISSN 0272-5428. LCCN TK7885.A1 S92 1978. IEEE order number 78 CH1397-9C.

IEEE:1979:ASF

[20] IEEE, editor. 20th Annual Symposium on Foundations of Computer Science: Oct. 29–31, 1979, San Juan, Puerto Rico. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1979. CODEN ASF-PDV. ISBN ???? ISSN 0272-5428. LCCN QA267 .S95 1979; TK7885.A1 S92 1979.

IEEE:1980:ASF

[21] IEEE, editor. 21st annual Symposium on Foundations of Computer Science: October 13–15, 1980, Syracuse, New York. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1980. CODEN ASF-PDV. ISBN ???? ISSN 0272-5428. LCCN QA76.6 .S95 1980; TK7885.A1 S92 1980. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog no. 80CH1498-5.

IEEE:1981:ASF

[22] IEEE, editor. 22nd Annual Symposium on Foundations of Computer Science: October 28–30, 1981, [Nashville, Tennessee: papers]. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1981. CODEN ASFPDV. ISBN ???? ISSN 0272-5428. LCCN TK7885.A1 S92 1981; QA76.6 .S95 1981. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog no. 81CH1695-6.

IEEE:1982:ASF

[23] IEEE, editor. 23rd annual Symposium on Foundations of Computer Science, November 3–5, 1982, Chicago, Illinois. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1982. CODEN ASF-PDV. ISBN ???? ISSN 0272-5428. LCCN QA76.6 .S95 1982. IEEE catalog no. 82CH1806-9. IEEE Computer Society order no. 440.

IEEE:1983:ASF

[24] IEEE, editor. 24rd Annual Symposium on Foundations of Computer Science: November 7–9, 1983, Tucson, Arizona. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1983. CODEN ASF-PDV. ISBN 0-8186-0508-1. ISSN 0272-5428. LCCN QA76.6 .S95 1983. IEEE catalog no. 83CH1938-0.

IEEE:1984:ASF

[25] IEEE, editor. 25th annual Symposium on Foundations of Computer Science, October 24–26, 1984, Singer Island, Florida. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1984. CO-DEN ASFPDV. ISBN 0-8186-8591-3, 0-8186-0591-X (paperback), 0-8186-4591-1 (microfiche). ISSN 0272-5428. LCCN QA 76 S979 1984. IEEE catalog no. 84CH2085-9.

IEEE:1985:ASF

[26] IEEE, editor. 26th annual Symposium on Foundations of Computer Science, October 21–23, 1985, Portland, Oregon. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1985. CODEN ASFPDV. ISBN 0-8186-4644-6 (microfiche), 0-8186-8644-8 (casebound), 0-8186-0644-4 (paperback). ISSN 0272-5428. LCCN QA 76 S979 1985. IEEE catalog no. 85CH2224-4. IEEE Computer Society order no. 644.

IEEE:1986:ASF

[27] IEEE, editor. 27th annual Symposium on Foundations of Computer Science, October 27–29, 1986, Toronto, ON, Canada. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1986. CODEN ASF-PDV. ISBN 0-8186-0740-8 (paperback), 0-8186-4740-X (microfiche), 0-8186-8740-1 (casebound). ISSN 0272-5428. LCCN QA 76 S979 1986. IEEE catalog no. 86CH2354-9. IEEE Computer Society order no. 740.

IEEE:1987:ASF

[28] IEEE, editor. 28th annual Symposium on Foundations of Computer Science, October 12–14, 1987, Los Angeles, California. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1987. CODEN ASF-PDV. ISBN 0-8186-0807-2, 0-8186-4807-4 (microfiche), 0-8186-8807-6 (casebound). ISSN 0272-5428. LCCN QA 76 S979 1987. IEEE Catalog no. 87CH2471-1. Computer Society order number 807.

IEEE:1988:ASF

[29] IEEE, editor. 29th annual Symposium on Foundations of Computer Science, October 24–26, 1988, White Plains, New York. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1988. CO-DEN ASFPDV. ISBN 0-8186-0877-3 (paperback), 0-8186-4877-5 (microfiche), 0-8186-8877-7 (hard). ISSN 0272-5428. LCCN QA 76 S979 1988. IEEE catalog no. 88CH2652-6. Computer Society order no. 877.

IEEE:1989:ASF

[30] IEEE, editor. 30th annual Symposium on Foundations of Computer Science, October 30–November 1, 1989, Research Triangle Park, North Carolina. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1989. CODEN ASFPDV. ISBN 0-8186-1982-1 (casebound), 0-8186-5982-3 (microfiche). ISSN 0272-5428. LCCN QA 76 S979 1989; TK7885.A1 S92 1989. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog no. 89CH2808-4. Computer Society order no. 1982.

IEEE:1990:PAS

[31] IEEE, editor. Proceedings: 31st Annual Symposium on Foundations of Computer Science: October 22–24, 1990, St. Louis, Missouri. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1990. CODEN ASFPDV. ISBN 0-8186-2082-X (paperback), 0-8186-6082-1 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1990. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog number 90CH29256. Computer Society order no. 2082.

IEEE:1991:PAS

[32] IEEE, editor. Proceedings: 32nd annual Symposium on Foundations of Computer Science, San Juan, Puerto Rico, October 1–4, 1991. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1991. CO-DEN ASFPDV. ISBN 0-8186-2445-0. ISSN 0272-5428. LCCN TK7885.A1 S92 1991. IEEE Catalog no. 91CH3062-7. Computer Society order no. 2445.

IEEE:1992:ASF

[33] IEEE, editor. 33rd Annual Symposium on Foundations of Computer Science: October 24–27, 1992, Pittsburgh, Pennsylvania: proceedings [papers]. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1992. CODEN ASFPDV. ISBN 0-8186-2901-0 (microfiche), 0-8186-2900-2 (paperback). ISSN 0272-5428. LCCN QA 76 S979 1992. IEEE Catalog Number 92CH3188-0. IEEE Computer Society Press Order Number 2900.

IEEE:1993:ASF

[34] IEEE, editor. 34th Annual Symposium on Foundations of Computer Science: November 3–5, 1993, Palo Alto, California: proceedings [papers]. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1993. CODEN ASFPDV. ISBN 0-8186-4370-6 (paperback), 0-8186-4371-4 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1993. IEEE Catalog Number 93CH3368-8. IEEE Computer Society Press Order Number 4372-02.

Goldwasser:1994:P

[35] Shafi Goldwasser, editor. Proceedings: 35th Annual Symposium on Foundations of Computer Science, November 20–22, 1994, Santa Fe, New Mexico. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1994. CODEN ASFPDV. ISBN 0-8186-6582-3. ISSN 0272-5428. LCCN QA 76 S979 1994. IEEE catalog number 94CH35717. IEEE Computer Society Press Order Number 6580-02.

IEEE:1995:ASF

[36] IEEE, editor. 36th Annual Symposium on Foundations of Computer Science: October 23–25, 1995, Milwaukee, Wisconsin. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1995. CODEN ASF-PDV. ISBN 0-7803-3121-4 (casebound), 0-8186-7183-1 (softbound), 0-8186-7184-X (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1995. IEEE catalog number 95CB35834.

IEEE:1996:ASF

[37] IEEE, editor. 37th Annual Symposium on Foundations of Computer Science: October 14–16, 1996, Burlington, Vermont. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1996. CODEN ASF-PDV. ISBN 0-7803-3762-X (casebound), 0-8186-7594-2 (softbound), 0-8186-7596-9 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1996. IEEE catalog number 96CH35973. IEEE Computer Society Press order number PR07594.

IEEE:1997:ASF

[38] IEEE, editor. 38th Annual Symposium on Foundations of Computer Science: October 20–22, 1997, Miami Beach, Florida. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1997. CODEN ASF-PDV. ISBN 0-8186-8197-7, 0-8186-8198-5 (casebound), 0-8186-8199-3 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 .S92 1997. IEEE catalog number 97CB36150. IEEE Computer Society Press order number PR08197.

IEEE:1998:ASF

[39] IEEE, editor. 39th Annual Symposium on Foundations of Computer Science: proceedings: November 8–11, 1998, Palo Alto, California. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1998. CODEN ASFPDV. ISBN 0-8186-9172-7 (softbound), 0-7803-5229-7 (casebound), 0-8186-9174-3 (microfiche). ISSN 0272-5428. LCCN QA267 .S95 1998 Sci-Eng. IEEE Catalog Number 98CB36280. IEEE Computer Society Press order number PR9172.

IEEE:1999:ASF

[40] IEEE, editor. 40th Annual Symposium on Foundations of Computer Science: October 17–19, 1999, New York City, New York. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1999. CO-DEN ASFPDV. ISBN 0-7695-0409-4 (softbound), 0-7803-5955-0 (casebound), 0-7695-0411-6 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1999. IEEE Catalog Number 99CB37039.

IEEE:2000:ASF

[41] IEEE, editor. 41st Annual Symposium on Foundations of Computer Science: proceedings: 12–14 November, 2000, Redondo Beach, California. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2000. CODEN ASFPDV. ISBN 0-7695-0850-2, 0-7695-0851-0 (case), 0-7695-0852-9 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 2000. IEEE Computer Society order number PR00850.

IEEE:2001:ISF

[42] IEEE, editor. 42nd IEEE Symposium on Foundations of Computer Science: proceedings: October 14–17, 2001, Las Vegas, Nevada, USA. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2001. CODEN ASFPDV. ISBN 0-7695-1390-5, 0-7695-1391-3 (case), 0-7695-1392-1 (microfiche). ISSN 0272-5428. LCCN ????

IEEE:2002:PAI

[43] IEEE, editor. Proceedings of the 43rd Annual IEEE Symposium on Foundations of Computer Science, FOCS 2002, Vancouver, BC, Canada, 16–19 November 2002. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2002. CODEN ASFPDV. ISBN 0-7695-1822-2. ISSN 0272-5428. LCCN

QA267.  URL http://ieeexplore.
ieee.org/iel5/8411/26517/01181875.

pdf. IEEE Computer Society order number PR01822.

IEEE:2003:PAI

[44] IEEE, editor. Proceedings: 44th Annual IEEE Symposium on Foundations of Computer Science, FOCS 2003, 11– 14 October 2003, Cambridge, Massachusetts. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2003. CODEN ASFPDV. ISBN 0-7695-2040-5. ISSN 0272-5428. LCCN QA76

.S979 2003. URL http://ieeexplore.
ieee.org/iel5/8767/27770/01238173.
pdf;   http://ieeexplore.ieee.org/
xpl/RecentCon.jsp?punumber=8767&

conhome=1000292. IEEE Computer Society order number PR02040.

IEEE:2004:PAI

[45] IEEE, editor. Proceedings: 45th Annual IEEE Symposium on Foundations of Computer Science: FOCS 2004, 17–19 October, 2004, Rome, Italy. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2004. CODEN ASFPDV. ISBN 0-7695-2228-9. ISSN 0272-5428. LCCN

QA276.  URL http://ieeexplore.
ieee.org/iel5/9430/29918/01366212.

                                        pdf;   http://ieeexplore.ieee.org/
                                        servlet/opac?punumber=9430. IEEE

Computer Society Order Number P2228.

IEEE:2005:AIS

[46] IEEE, editor. 46th Annual IEEE Symposium on Foundations of Computer Science: FOCS 2005: 23–25 October, 2005, Pittsburgh, Pennsylvania, USA. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2005. ISBN 0-7695-2468-0, 0-7695-2362-5. ISSN 0272-5428. LCCN QA76 .S979 2005. IEEE Computer Society order number P2468.

IEEE:2006:AIS

[47] IEEE, editor. 47th Annual IEEE Symposium on Foundations of Computer Science: FOCS 2006: 21–24 October, 2006, Berkeley, California. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2006. ISBN 0-7695-2720-5, 0-7695-2362-5. ISSN 0272-5428. LCCN QA76 .S974

                                        2006. URL http://ieeexplore.ieee.
                                        org/servlet/opac?punumber=4031329.

IEEE Computer Society Order Number P2720.

IEEE:2007:PAI

[48] IEEE, editor. Proceedings of the 48th Annual IEEE Symposium on Foundations of Computer Science: [FOCS 2007]: October 20–23, 2007, Providence, Rhode Island. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2007. ISBN 0-7695-3010-9. ISSN 0272-5428. LCCN QA76 .S974 2007. URL http:/

                                        /ieeexplore.ieee.org/servlet/opac?
                                        punumber=4389466. IEEE Computer So-

ciety order number P3010.

IEEE:2008:PAI

[49] IEEE, editor. Proceedings of the 49th Annual IEEE Symposium on Foundations of Computer Science: October 25–23, 2008, Philadelphia, Pennsylvania, USA. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2008. ISBN 0-7695-3436-8. ISSN 0272-5428. LCCN QA76 .S95

2008. URL http://ieeexplore.ieee.
org/servlet/opac?punumber=4690923.

IEEE Computer Society order number P3436.

IEEE:2009:PAI

[50] IEEE, editor. Proceedings of the 50th Annual IEEE Symposium on Foundations of Computer Science: October 25–27, 2009, Atlanta, Georgia, USA. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2009. ISBN 0-7695-3850-9. ISSN 0272-5428. LCCN QA76 .S95 2009. IEEE Computer Society order number P3850.

IEEE:2010:PIA

[51] IEEE, editor. Proceedings of the 2010 IEEE 51st Annual Symposium on Foundations of Computer Science: 23–26 October 2010, Las Vegas, Nevada, USA. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2010. ISBN 0-7695-4244-1, 1-4244-8525-8. ISSN 0272-5428. LCCN QA76 .S95 2010. URL http://

opac.ieeecomputersociety.org/opac?
year=2010&volume=00&catalog=4244&

acronym=focs. IEEE Computer Society order number P4244.

IEEE:2011:PIA

[52] IEEE, editor. Proceedings of the 2011 IEEE 52nd Annual Symposium on Foundations of Computer Science: 22– 25 October 2011, Palm Springs, California, USA. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2011. ISBN 0-7695-4571-8, 1-4577-1843-X. ISSN 0272-5428. LCCN QA76 .S95 2011. URL http://

opac.ieeecomputersociety.org/opac?
year=2011&volume=00&catalog=4571&

acronym=focs. IEEE Computer Society order number P4571.

IEEE:2012:PIA

[53] IEEE, editor. Proceedings of the 2012 IEEE 53rd Annual Symposium on Foundations of Computer Science: 20–23 October 2012, Hyatt Regency, New Brunswick, New Jersey, USA. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2012. ISBN 1-4673-4383-8. ISSN 0272-5428. LCCN QA76 .S95 2012. URL

http://dimacs.rutgers.edu/FOCS12/;
 http://theory.stanford.edu/~tim/

focs12/. IEEE Computer Society order number P????.

IEEE:2013:PIA

[54] IEEE, editor. Proceedings of the 2012 IEEE 54th Annual Symposium on Foundations of Computer Science: 26–29 October 2013, Berkeley, CA, USA. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2013. ISBN 0-7695-5135-1. ISSN 0272-5428. LCCN QA76 .S95 2012. IEEE Computer Society order number P????.